(3.3.13) The tallest living man at one time had a height of 244 cm. The shortest living man at that time had a height of 65.5 cm. Heights of men at that time had a mean of 172.29 cm and a standard deviation of 7.64 cm. Which of these two men had the height that was more​ extreme?

Since the z score for the tallest man is z= and the z score for the shortest man is z=
​, the tallest/shortest man had the height that was more extreme (Round to two decimals)
1
Expert's answer
2020-11-23T18:30:16-0500

Z score for the tallest person is "\frac{244-172.29}{7.64}=9.39"

The z score for the shortest person is "\frac{65.5-172.29}{7.64}=-13.98"

Since the z score for the tallest man is z=9.39 and the z score for the shortest man is z=-13.98, the shortest man had the height that was more extreme.